OpenStudy (anonymous):

how to get the derivative of 4-x^2?

OpenStudy (abb0t):

\[\frac{ d }{ dx }a^n = na^{n-1}\] where n is some constant.

OpenStudy (zehanz):

If you may use the Power Rule: \(\left(x^n\right)'=nx^{n-1}\), this is not that hard. YOu then only have to know that the derivative of a constant (4) is 0. The "-" sign just get carried over.

OpenStudy (zehanz):

On the other hand, it could be that you have to find the derivative using the limit-definition. Could you tell me if that is the case?

OpenStudy (zehanz):

@abb0t: you'd better change that a to x, or x to a :)

OpenStudy (anonymous):

ok I know that 2x^2-1=2x?

OpenStudy (zehanz):

You mean: \(\left(x^2\right)'=2x^{2-1}=2x^1=2x\), yes, that is right, so what is the final answer?

OpenStudy (anonymous):

yes, that's what I meant... for the final answer is what I am stuck at?

OpenStudy (zehanz):

It is the minus-sign: \(\left(4-x^2\right)'=0-2x=-2x\)

OpenStudy (anonymous):

oh ok I see thanks @ZeHanz

OpenStudy (zehanz):

YW!

OpenStudy (amistre64):

i would find the derivative by taking the limit as h tends to zero of the difference quotient\[\lim_{h\to~0}\frac{(4-(x+h)^2)-(4-x^2)}{x+h-x}\]